OpenStudy (anonymous):

What is the equation of the line that passes through the point (0, 3) and is perpendicular to the line y = 2x + 3?

OpenStudy (anonymous):

negative reciprocal of the slope is the perpendicular slope

OpenStudy (anonymous):

first of all find the slope of y=2x +3 .... it's 2, this means the perpendicular slope is -1/2 We now got A in y=Ax+b Going through point 0,3 ... fill in formula; find b=3

OpenStudy (anonymous):

so would it be y=-1/2x+3??

OpenStudy (mathmate):

Right, assuming the x goes with the numerator, y=-(1/2)x+3